# Vector Quantization using Gaussian Variational Autoencoder

### v1

@@ Appendix B Stable Mean-KL Parametrization
We investigate an alternative to TDC, namely the Mean-KL parametrization (Lin2023MinimalRC), which is considered to be easier to train than TDC since it does not require the construction of an empirical $R(\lambda)$ model.
@@ B.1 Mean-KL Parametrization
The Mean-KL parametrization (Lin2023MinimalRC) supports grouping with $m>1$. Its neural network output consists of two $m$-dimensional tensors, $\hat{\gamma}{i:i+m}$ and $\tau{i:i+m}$, which allocate the $D_{\mathrm{KL}}$ target $K$ across the $m$ dimensions and determine the mean, respectively. More specifically, the Mean-KL parametrization determines the mean $\mu_{i:i+m}$ and variance $\sigma^{2}_{i:i+m}$ as follows, where $\mathcal{W}(\cdot)$ denotes the principal branch of the Lambert W function:
@@ B.3 Results of Stable Mean-KL Parametrization
Table 9: Quantitive results on ImageNet validation dataset.
Method
bpp
(# of tokens)
UNet based ViT based
@@ C.1 Details of Training and Distortion Objective
We train all VQ-VAEs on the ImageNet validation dataset using $8\times$ H100 GPUs for approximately 24 hours. For UNet models, we train each model for 30 epochs using the ADAM (Kingma2014AdamAM) optimizer with a batch size of $128$ and a learning rate of $1\times 10^{-4}$. For ViT models, we train each model for 40 epochs using the ADAM optimizer with a batch size of $256$ and a learning rate of $4\times 10^{-7}$.
All VQ-VAEs are trained using the following distortion objective, which corresponds to the classical VQ-GAN (Esser2020TamingTF) objective employed in the Stable Diffusion VAE (rombach2022high).
@@
Following the implementation of Stable Diffusion, we set $w_{1}=1.0$ and $w_{2}=0.75$ for UNet models. Consistent with the implementation of BSQ (zhao2024image), we set $w_{1}=0.1$ and $w_{2}=0.1$ for ViT models.
@@
For the image generation model, we first train all VQ-VAEs using images of size $128\times 128$, following the same settings as described above. Subsequently, we train the auto-regressive transformer for image generation using the implementation of IBQ (Shi2024TamingSV) with a Llama-base transformer architecture. The transformer has a vocabulary size of $2^{16}$, $16$ layers, $16$ attention heads, and an embedding dimension of $1024$. We train the transformer for $100$ epochs using the ADAM optimizer with a learning rate of $3\times 10^{-4}$ and a batch size of $512$.
C.2 Details of Grouping Strategies
We extend the notation from main text. We group $m$ tokens into one large token with codebook size $K$. For each quantization output $\hat{z}_{i}$, we denote the corresponding index in the codebook as $\mathcal{I}_{i}$. And we have $c_{\mathcal{I}_{i}}=\hat{z}_{i}$.
Post Quantization (PQ) The Post Quantization (PQ) grouping strategy happens after GQ. We first train a Gaussian VAE with codebook size for each dimension $K^{1/m}$. Next, we obtain GQ tokens $\mathcal{I}_{1:d}$. Then, we group those $d$ tokens into $d/m$ groups with group size of $m$. Denote the group index as $g=0,\ldots,d/m-1$, then each group can be denoted as $\{\mathcal{I}_{gm+l}\},l=1,\ldots,m$.
In that case, we have $\max\{\mathcal{I}_{gm+l}\}\leq K^{1/m}$. Then, we can view each $\mathcal{I}_{gm+l}$ as an integer in a $K^{1/m}$-base numerical system. Then, aggregating $m$ tokens $\{\mathcal{I}_{gm+l}\}$ into one large token $I^{m}_{g}$ is as easy as concatenating $m$ tokens into a larger integer in a $K$-based numerical system: $\mathcal{I}_{g}^{m}=\sum_{l=1}^{m}\mathcal{I}_{gm+l}K^{(l-i)/m}\leq K$.
Post Training (PT) The Post Training (PT) grouping strategy happens after the training of Gaussian VAE. We still start with a Gaussian VAE with target codebook size for each dimension $K^{1/m}$. Next, instead of performing GQ for each dimension, we consider the following $m$ dimension GQ for each group $g$:
[…]
In fact, we can view one dimension GQ as a max likelihood:
$\hat{z}_{i}=\arg\max_{c_{j}\in\{c_{1:K}\}}\log q(Z_{i}=c_{j}|X),\textrm{ where }c_{1:K}\sim\mathcal{N}(0,1).$
Then we can extend the max likelihood to $m$ dimension, which is equivalent to the basic version of PT in Eq. 32.
$\hat{z}_{gm:gm+m}=\arg\max_{c_{j}\in\{c_{1:K}\}}\log q(Z_{gm:gm+m}=c_{j}|X),\textrm{ where }c_{1:K}\sim\mathcal{N}(0,I_{m}).$
When group size $m$ is large, we observe that vanilla PT in Eq. 32 leads to codebook collapse (See Table. 13) and a decay in performance. Therefore, we include a regularization term weighted by $\omega$:
[…]
Training Aware (TA) The Training Aware (TR) grouping strategy happens before the training of Gaussian VAE. More specifically, we directly interoperate grouping with the TDC, and introduce the $m$ group TDC as follows:
$\mathcal{L}_{TDC}^{m}=\sum_{g=0}^{d/m}\lambda_{g}\sum_{l=1}^{m}D_{KL}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))+\Delta(X,g(z)),$
$\textup{where }\lambda_{g}=\left\{\begin{array}[]{ll}\lambda_{\textrm{min}},&\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))<{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K-\alpha}\textup{ bits},\\ \lambda_{\textrm{mean}},&\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))\in[\log_{2}K-0.5,\log_{2}K+0.5]\textup{ bits},\\ \lambda_{\textrm{max}},&\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K+\alpha}\textup{ bits}.\end{array}\right$. (36)
Besides, the $\lambda$ update heuristic should also consider $m$ dimension as a group:
$\lambda_{\textrm{min}}=\lambda_{\textrm{min}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ min}_{g}\{\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))\}>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K-\alpha}\textrm{ else }\lambda_{\textrm{min}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta},$
$\lambda_{\textrm{mean}}=\lambda_{\textrm{mean}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ mean}_{g}\{\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))\}>\log_{2}K\textrm{ else }\lambda_{\textrm{mean}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta},$
$\lambda_{\textrm{max}}=\lambda_{\textrm{max}}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}\textrm{ if}\textrm{ max}_{g}\{\sum_{l=1}^{m}D_{KL(2)}(q(Z_{gm+l}|X)||\mathcal{N}(0,1))\}>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{2}K+\alpha}\textrm{ else }\lambda_{\textrm{max}}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}.$ (37)
C.3 Details of Hyper-parameters
@@
VQGAN (van2017neural) We adopt the factorized codebook VQGAN variant following (Zheng2022MoVQMQ). For each codebook, we use a codebook size of $K=2^{16}$ and a group dimension of $m=16$. The number of codebooks $n$ varies depending on the bitrate. Additionally, we use a codebook loss weight of $\lambda=1.0$ and a commitment loss weight of $\zeta=0.25$.
FSQ (mentzer2023finite) The only parameter of FSQ is the codebook list $l$, which represents the quantization level for each integer value. We set each unit value to $2^{4}=16$, and populate $l$ with the appropriate number of $16$s according to the desired bitrate.
LFQ (yu2023language) For LFQ at $0.25$ bpp, we follow the original paper and split a large codebook of size $2^{16}$ into $n=2$ smaller codebooks, each with $K=2^{8}$ and a codebook dimension of $m=8$. We use an entropy loss weight of $\lambda=0.1$ and a commitment loss weight of $\zeta=0.025$.
BSQ (zhao2024image) We fix the size of each BSQ codebook to $2^{1}$, with a group dimension of $m=1$, and vary the number of codebooks $n$ according to the desired bitrate. For the entropy penalization parameter, we set $\lambda=0.1$, following the official implementation.
GQ We use TR grouping with a fixed codebook size of $K=2^{16}$. Each group has dimension $m$, and there are $n$ groups in total, such that $m\times n=16$. The value of $m$ varies depending on the bitrate.
@@ 1.00 $l=\{16,16,16,16,16,16,16,16,16,16,16,16,16,16,16,16\}$
Appendix D Additional Quantitive Results
D.1 Comparison to other Conversion Methods
In Table. 11, we compare our GQ to TokenBridge (wang2025bridging) and ReVQ (zhang2025quantizethenrectifyefficientvqvaetraining), which also convert a Gaussian VAE into a VQ-VAE. Those results are from their original paper. It is shown that GQ has best reconstruction metrics. Besides, only GQ has theoretical guarantee.
Table 11: Quantitive results on ImageNet dataset. Bold: best, : from paper, -: not available.
Method
Training
free
Theoretical
guarantee
bpp
(# of tokens)
[…]
OpenMagViT-V2 No No 0.07 ($2^{18}\times 256$) 21.63 0.111 0.640 1.17
TokenBridge Yes No 0.375 ($2^{6}\times 4096$) - - - 1.11
ReVQ-256T No No 0.07 ($2^{18}\times 256$) 21.96 0.121 0.640 2.05
GQ (Ours) Yes Yes 0.07 ($2^{18}\times 256$) 22.30 0.116 0.642 1.04
D.2 The TDC Parameters
In Table 12, we show the effect of different TDC parameters. It is shown that $\beta=1.001,1.01,1.1$ does not make a significant different on result. However, setting $\alpha>0.5$ is harmful to performance.
Table 12: Ablation study on TDC parameters.
[…]
0.5 1.01 27.61 0.059 0.807 0.529
0.1 1.01 27.56 0.058 0.812 0.551
1.0 1.01 27.61 0.063 0.811 0.701
0.5 1.1 27.63 0.060 0.809 0.534
0.5 1.001 27.48 0.058 0.804 0.510
D.3 Effectiveness of Grouping Strategies
In Table 13, we evaluate the effect of token grouping techniques. The scenario we consider involves grouping four $4$-bit tokens into a single $16$-bit token, which is a reasonable setting for autoregressive generation. The results show that PQ has no effect on reconstruction performance, while PT provides some improvements in PSNR and SSIM. In contrast, TR, which involves training the Gaussian VAE with a grouping target, achieves the best reconstruction performance.
Table 13: Effects of token grouping on ImageNet dataset. TR strategy has best reconstruction performance for grouping $m=4$.
[…]
Gaussian VAE (w/ TDC) no (m=1) - 4.26, 2.93-5.63 1.06 32.61 0.023 0.906 0.460
GQ no (m=1) 4 - 1.00 32.11 0.023 0.906 0.414
GQ PQ (m=4) 16 - 1.00 32.11 0.023 0.906 0.414
GQ PT (m=4) 16 - 1.00 32.15 0.023 0.907 0.428
Gaussian VAE (w/ TDC) TR (m=4) - 15.99, 14.81-17.54 1.00 32.62 0.023 0.909 0.331
GQ TR (m=4) 16 - 1.00 32.47 0.023 0.907 0.322
Additionally, in Table 14, we show the effect of the regularization parameter $\omega$ for PT and TR. For high bitrates, such as $1.00$ bpp, regularization is not required; in other words, setting $\omega=0.0$ yields the good enough codebook usage and rFID. For lower bitrates, such as $0.50$ bpp, $\omega=0.0$ leads to codebook collapse, while $\omega=2.0$ achieves the best codebook entropy and rFID.
@@ D.3 Effectiveness of Grouping Strategies
Table 15: Quantitive results on COCO 2017 dataset. Bold: best.
Method
bpp
(# of tokens)
UNet based ViT based
@@
VQGAN
0.50
($^{\textup{16}}\times2$048)
29.06 0.049 0.839 6.616 27.83 0.058 0.832 5.461
@@
D.4 The Quantization Error in Pixel Space
@@
Corollary 3. Following the setting in Theorem 1, and assuming the decoder $g(.)$ satisfy $|g(x_{1})-g(x_{2})|\leq c_{3}|x_{1}-x_{2}|$, we have:
$Pr\{|g(\hat{z}_{i})-g(\mu_{i})|\geq c_{3}\sigma_{i}\}\leq\exp(-e^{t}\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5}).$ (38)
Proof. As $|g(\hat{z}_{i})-g(\mu_{i})|\leq c_{3}|\hat{z}_{i}-\mu_{i}|$, we have $Pr\{|g(\hat{z}_{i})-g(\mu_{i})|\geq c_{3}\sigma_{i}\}\leq Pr\{c_{3}|\hat{z_{i}}-\mu_{i}|\geq c_{3}\sigma_{i}\}=Pr\{|\hat{z_{i}}-\mu_{i}|\geq\sigma_{i}\}$.
@@
D.5 Robustness of Codebook to Random Seed
As the codebook is usually large enough ($2^{16}$) to compensate for the randomness. We provide additional in Table 17 showing that the random seed has little effect on reconstruction performance. We used three continuos random seeds without cherry-picking. It is clear that the performance of GQ is not affected by randomness.
Table 17: The effect of codebook randomness on the performance of GQ.
[…]
42 27.61 0.059 0.807 0.529
43 27.61 0.059 0.807 0.523
44 27.62 0.059 0.807 0.526
D.6 The Effect of Simply Increasing Codebook Size
According to Theorem 1, increasing $\log K$ too much over $D_{KL}$ will not significantly improve reconstruction but will lead to a waste of bitrate (tokens). We provide an additional in Table 18, showing that GQ trained with $14$ bits and quantized into $18$ bits does not perform as well as GQ trained with $18$ bits and quantized into $18$ bits. The drawback of simply increasing the codebook size for GQ is that it is not as effective as increasing the target of TDC to that size and quantizing with a proper $\log K=D_{KL}$.
Table 18: The effect of simply increase codebook size.
[…]
$D_{KL}(2)=14$ $\log_{2}K=14$ $2^{14}\times 1024$ 25.31 0.064 0.762 0.491
$D_{KL}(2)=14$ $\log_{2}K=18$ $2^{18}\times 1024$ 27.79 0.059 0.808 0.513
$D_{KL}(2)=18$ $\log_{2}K=18$ $2^{18}\times 1024$ 27.86 0.054 0.804 0.424
D.7 Quantized Latent Visualization
In Figure 4, we show the t-NSE (Maaten2008VisualizingDU) visualization of latent after GQ, using 5 subclass of ImageNet.
@@ D.7 Quantized Latent Visualization
D.8 More Generation Results
@@
D.9 Prior-Posterior Mismatch
@@
$D_{KL}(q(Z)||N(0,1))\approx\frac{1}{N}\sum_{i=1}^{N}(\log q(z^{i})-\log N(z^{i}|0,I)).$ (39)
@@
We train a diffusion model to estimate $\log q(z^{i})$ by using PF-ODE and Skilling-Hutchinson trace estimator (See Appendix D.2 of Song2020ScoreBasedGM). We use the Gaussian VAE (w/ TDC) + DiT diffusion model and ImageNet validation dataset. The euler PF-ODE steps is set to 250 and the Skilling-Hutchinson number of sample is set to 1. The result is shown in Table 21. The results show that the prior-posterior mismatch is only 0.00033 bits per pixel, accounting for approximately 0.1% of the total bpp. Furthermore, the best bpp and the actual bpp show no significant difference on a scale of 0.01. This indicates that the ”bitrate waste” caused by the prior-posterior mismatch is negligible, and the mismatch itself is not significant.
@@
Divergence bits per pixel
@@
prior posterior mismatch ($D_{KL}(q(Z)\|\|N(0,1))$) 0.000328
D.10 Complexity

### v2

@@ Appendix B Stable Mean-KL Parametrization
We investigate an alternative to TDC, namely the Mean-KL parametrization (Lin et al., 2023), which is considered to be easier to train than TDC since it does not require the construction of an empirical $R(\lambda)$ model.
@@ B.1 Mean-KL Parametrization
The Mean-KL parametrization (Lin et al., 2023) supports codebook dimension $m>1$. Its neural network output consists of two $m$-dimensional tensors, $\hat{\gamma}_{i:i+m}$ and $\tau_{i:i+m}$, which allocate the $R_{i:i+m}$ target $\log K$ across the $m$ dimensions and determine the mean, respectively. More specifically, the Mean-KL parametrization determines the mean $\mu_{i:i+m}$ and variance $\sigma^{2}_{i:i+m}$ as follows, where $\mathcal{W}(\cdot)$ denotes the principal branch of the Lambert $\mathcal{W}$ function:
@@ B.3 Results of Stable Mean-KL Parametrization
Table 12: Quantitative results on ImageNet validation dataset comparing GQ with different constraint.
Method bpp ($K\times N$) UNet based ViT based
@@ C.1 Details of Training and Distortion Objective
We train all VQ-VAEs on the ImageNet validation dataset using $8\times$ H100 GPUs for approximately 24 hours. For UNet models, we train each model for 30 epochs using the Adam (Kingma and Ba, 2014) optimizer with a batch size of $128$ and a learning rate of $1\times 10^{-4}$. For ViT models, we train each model for 40 epochs using the Adam optimizer with a batch size of $256$ and a learning rate of $4\times 10^{-7}$.
All VQ-VAEs are trained using the following distortion objective, which corresponds to the classical VQ-GAN (Esser et al., 2020) objective employed in the Stable Diffusion VAE (Rombach et al., 2022).
@@
Following the implementation of Stable Diffusion, we set $w_{1}=1.0$ and $w_{2}=0.75$ for UNet models. Consistent with the implementation of BSQ (Zhao et al., 2024), we set $w_{1}=0.1$ and $w_{2}=0.1$ for ViT models.
@@
For the image generation model, we first train all VQ-VAEs using images of size $128\times 128$, following the same settings as described above. Subsequently, we train the auto-regressive transformer for image generation using the implementation of IBQ (Shi et al., 2024b) with a Llama-base transformer architecture. The transformer has a vocabulary size of $2^{16}$, $16$ layers, $16$ attention heads, and an embedding dimension of $1024$. We train the transformer for $100$ epochs using the Adam optimizer with a learning rate of $3\times 10^{-4}$ and a batch size of $512$.
C.2 Details of Hyper-parameters
@@
VQGAN (Van Den Oord et al., 2017) We adopt the factorized codebook VQGAN variant following (Zheng et al., 2022). For each codebook, we use a codebook size of $K=2^{16}$ and a group dimension of $m=16$. The number of codebooks $n$ varies depending on the bitrate. Additionally, we use a codebook loss weight of $\lambda=1.0$ and a commitment loss weight of $\zeta=0.25$.
FSQ (Mentzer et al., 2023) The only parameter of FSQ is the codebook list $l$, which represents the quantization level for each integer value. We set each unit value to $2^{4}=16$, and populate $l$ with the appropriate number of $16$s according to the desired bitrate.
LFQ (Yu et al., 2023) For LFQ at $0.25$ bpp, we follow the original paper and split a large codebook of size $2^{16}$ into $n=2$ smaller codebooks, each with $K=2^{8}$ and a codebook dimension of $m=8$. We use an entropy loss weight of $\lambda=0.1$ and a commitment loss weight of $\zeta=0.025$.
BSQ (Zhao et al., 2024) We fix the size of each BSQ codebook to $2^{1}$, with a group dimension of $m=1$, and vary the number of codebooks $n$ according to the desired bitrate. For the entropy penalization parameter, we set $\lambda=0.1$, following the official implementation.
GQ We use a fixed codebook size of $K=2^{16}$. We adopt fixed $\alpha=0.5,\beta=1.01$. We adopt $m=16,8,4,\omega=2.0,2.0,0.0$ for bpp=0.25,0.5,1.0=0.25,0.5,1.0 respectively.
@@ 1.00 $l=\{16,16,16,16,16,16,16,16,16,16,16,16,16,16,16,16\}$
Appendix D Additional Quantitative Results
D.1 Detail in Codebook Usage Hyper-parameter $\omega$
Additionally, in Table 14, we show the effect of the codebook usage regularization parameter $\omega$. For high bitrates, such as $1.00$ bpp, regularization is not required; in other words, setting $\omega=0.0$ yields the good enough codebook usage and rFID. For lower bitrates, such as $0.50$ bpp, $\omega=0.0$ leads to codebook collapse, while $\omega=2.0$ achieves the best codebook entropy and rFID.
@@ D.1 Detail in Codebook Usage Hyper-parameter $\omega$
Table 15: Quantitative results on COCO 2017 dataset. Bold: best.
Method bpp ($K\times N$) UNet based ViT based
@@
VQGAN 0.50 ($^{\textup{16}}\times2$048) 29.06 0.049 0.839 6.616 27.83 0.058 0.832 5.461
@@
D.2 The Quantization Error in Pixel Space
@@
Corollary 3. Following the setting in Theorem 3.1, and assuming the decoder $g(.)$ satisfy $|g(x_{1})-g(x_{2})|\leq c_{3}|x_{1}-x_{2}|$, we have:
$\textup{Pr}\{|g(\hat{z})-g(\mu)|\geq c_{3}\sigma_{i}\}\leq\exp(-e^{t}\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5}).$ (39)
Proof. As $|g(\hat{z}_{i})-g(\mu_{i})|\leq c_{3}|\hat{z}_{i}-\mu_{i}|$, we have $\textrm{Pr}\{|g(\hat{z}_{i})-g(\mu_{i})|\geq c_{3}\sigma_{i}\}\leq\textrm{Pr}\{c_{3}|\hat{z_{i}}-\mu_{i}|\geq c_{3}\sigma_{i}\}=\textrm{Pr}\{|\hat{z_{i}}-\mu_{i}|\geq\sigma_{i}\}$.
@@
D.3 Quantized Latent Visualization
In Figure 3, we show the t-NSE (van der Maaten and Hinton, 2008) visualization of latent after GQ, using 5 subclass of ImageNet.
@@ D.3 Quantized Latent Visualization
D.4 More Generation Results
@@
D.5 Prior-Posterior Mismatch
@@
$D_{KL}(q(Z)||\mathcal{N}(0,1))\approx\frac{1}{N}\sum_{i=1}^{N}(\log q(z^{i})-\log\mathcal{N}(z^{i}|0,I)).$ (40)
@@
We train a diffusion model to estimate $\log q(z^{i})$ by using PF-ODE and Skilling-Hutchinson trace estimator (See Appendix D.2 of Song et al. (2020)). We use the Gaussian VAE (w/ TDC) + DiT diffusion model and ImageNet validation dataset. The euler PF-ODE steps is set to 250 and the Skilling-Hutchinson number of sample is set to 1. The result is shown in Table 19. The results show that the prior-posterior mismatch is only 0.00033 bits-per-pixel, accounting for approximately 0.1% of the total bpp. Furthermore, the best bpp and the actual bpp show no significant difference on a scale of 0.01. This indicates that the ”bitrate waste” caused by the prior-posterior mismatch is negligible, and the mismatch itself is not significant.
@@
Divergence bits-per-pixel
@@
prior posterior mismatch ($D_{KL}(q(Z)\|\|\mathcal{N}(0,1))$) 0.000328
D.6 Complexity